# Finding all solutions of qKZ equations in characteristic $p$

### v1

@@ 4.3. Step $-\kappa$
(4.9) $Q^{\ell p-1}(z;-\kappa),\qquad\ell=1,\dots,\left[\frac{n(p-k)}{p}\right],$
are solutions of the qKZequations with step $-\kappa$ and values in $V$.
If $p$ does not divide $n$, then the total number of $p$-hypergeometric solutions of the qKZ equations with values in $V$ and steps $\kappa$ and $-\kappa$ equals
(4.10) $\left[\frac{n{k}}{p}\right]+\left[\frac{n(p-{k})}{p}\right]=n-1=\dim V.$
4.4. $p$-hypergeometric solutions of KZ equations
In this section, we remind the construction in [SV] of polynomial solutions modulo $p$ of the differential KZ equations with values in $V={\operatorname{Sing}}L^{\otimes n}[n-2]$.
Let $0<k<p$ be the positive integers such that $\kappa k\,\equiv\,-1\pmod{p}.$ Define master polynomial
$\bar{\Phi}(t,z,\kappa)=\prod_{a=1}^{n}(t-z_{a})^{k}\,.$
For $1\leqslant a\leqslant n$, define the weight functions
@@ 4.4. $p$-hypergeometric solutions of KZ equations
(4.11) $\bar{Q}(t,z;\kappa)=(\bar{Q}_{1}(t,z;\kappa),\dots,\bar{Q}_{n}(t,z;\kappa))^{\intercal}=\sum_{i}\bar{Q}^{i}(z;\kappa)t^{i}\,,$
where
$\bar{Q}^{i}(z;\kappa)=(\bar{Q}^{i}_{1}(z;\kappa),\dots,\bar{Q}^{i}_{n}(z;\kappa))^{\intercal}$
are vectors of polynomials in $z$.
Theorem 4.5 ([SV]).
@@
(4.12) $\bar{Q}^{\ell p-1}(z,\kappa),\qquad\ell=1,\dots,\left[\frac{n{k}}{p}\right],$
@@
Notice that $\bar{Q}_{a}(t,z;\kappa)$ are homogeneous polynomials in variables $t,z$ of degree $nk-1$, and $\bar{Q}^{\ell p-1}(z;\kappa)$ are vectors of homogeneous polynomials in $z$ of degree $nk-\ell p$.
4.5. Top-degree part of $p$-hypergeometric solutions
@@
(4.13) $P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}a_{d_{0},\dots,d_{n+1}}\,t^{d_{0}}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\alpha^{d_{n+1}}\,.$
@@
(4.14) $P(t,z_{1},\dots,z_{n},0)=\sum_{d_{0}+\dots+d_{n}=d}a_{d_{0},\dots,d_{n},0}\,t^{d_{0}}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\,.$
@@
$\bar{P}_{e}(z_{1},\dots,z_{n})=\sum_{e+d_{1}+\dots+d_{n}=d}a_{e,d_{1},\dots,d_{n},0}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\,,$
@@
(4.15) $P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}b_{d_{0},\dots,d_{n+1}}\,(t;\alpha)_{d_{0}}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\alpha^{d_{n+1}}\,.$
@@
$P_{e}(z_{1},\dots,z_{n},\alpha)=\sum_{e+d_{1}+\dots+d_{n+1}=d}b_{e,d_{1}+\dots,d_{n+1}}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\alpha^{d_{n+1}}\,,$
@@
(4.16) $P_{e}(z_{1},\dots,z_{n},0)=\sum_{e+d_{1}+\dots+d_{n}=d}b_{e,d_{1}+\dots,d_{n},0}\,z_{1}^{d_{1}}\dots z_{n}^{d_{n}}\,.$
@@
By formula (3.6), we have
(4.17) $a_{e,d_{1},\dots,d_{n},0}=b_{e,d_{1},\dots,d_{n},0}$
for any $e,d_{1},\dots,d_{n}$. This implies the lemma. ∎
Corollary 4.7 ([MV]).
Let $Q^{\ell p-1}(z;\kappa)$ be a $p$-hypergeometric solution of the qKZ equations from Theorem 4.3. Then
$Q^{\ell p-1}(z;\kappa)=\bar{Q}^{\ell p-1}(z;\kappa)+\dots,$
where $\bar{Q}^{\ell p-1}(z;\kappa)$ is the corresponding $p$-hypergeometric solution of the KZ equations from Theorem 4.5, and the dots denote the terms of degree less than $nk-\ell p=\deg\bar{Q}^{\ell p-1}(z;\kappa)$.
@@
5. Linear independence
5.1. Lexicographical order
Define length-lexicographical order on monomials $v=\\ z_{1}^{d_{1}}\dots z_{n}^{d_{n}}$: $v<w$ if degree$(v)<$degree$(w)$ or if degree$(v)=$degree$(w)$ and $v$ is smaller than $w$ in lexicographical order.
For a polynomial $f(z)=$ $\sum_{d_{1},\dots,d_{n}}a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\dots z_{n}^{d_{n}}$  denote by $\text{L}f(z)$ the nonzero summand $a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\dots z_{n}^{d_{n}}$ with the lexicographically largest monomial $z_{1}^{d_{1}}\dots z_{n}^{d_{n}}$. We call $\text{L}f(z)$ the leading term of $f(z)$.
5.2. Leading terms
@@
We have $r(1)>r(2)>\dots$. Denote
$a$ $=$ $(n-r(\ell))k-\ell p,$
(5.1) $u_{\ell}$ $=$ $\frac{(-1)^{nk-\ell p}}{k}\binom{k}{a}\left(0,\dots,0,k-a,k,\dots,k\right)^{\intercal}\ \in{\mathbb{F}}_{p}^{n}\,,$
where $k-a$ is the $r(\ell)+1$-st coordinate. Notice that the integers $k$ and $k-a$ are not divisible by $p$, and $u_{\ell}\in{\mathbb{F}}_{p}^{n}$ is a singular vector since the sum of its coordinates equals zero.
Lemma 5.1 ([V, Lemma 3.1]).
Assume that $p$ does not divide $n$ and $\ell\in\left\{1,\dots,d(\kappa)\right\}$. Then the leading term of $\bar{Q}^{\ell p-1}(z;\kappa)$ is given by the formula
$\operatorname{L}\!\bar{Q}^{\ell p-1}(z;\kappa)=(z_{1}\dots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}\,u_{\ell}\,.$
@@
(5.2) $\operatorname{L}\!Q^{\ell p-1}(z;\kappa)=\operatorname{L}\!\bar{Q}^{\ell p-1}(z;\kappa).$
@@
(5.3) $\operatorname{L}\!Q^{\ell p-1}(z;\kappa)=(z_{1}\dots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}\,u_{\ell}\,.$
@@
Let $n=3$ and $d(\kappa)=\left[\frac{3k}{p}\right]=1$. Then
(5.4) $\frac{p}{3}<k<\frac{2p}{3}\,.$
In this case there is exactly one $p$-hypergeometric solution $Q^{p-1}(z;\kappa)$. If $\frac{p}{2}<k<\frac{2p}{3}$, then the leading term of $Q^{p-1}(z;\kappa)$ is
(5.5) $z_{1}^{k}z_{2}^{2k-p}\frac{(-1)^{3k-p}}{k}\binom{k}{2k-p}\left(0,p-k,k\right)^{\intercal}\,.$
@@
(5.6) $z_{1}^{3k-p}\frac{(-1)^{k}}{k}\binom{k}{3k-p}\left(p-2k,k,k\right)^{\intercal}\,.$
Consider the collections of the $n$-vectors $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, as an $n\times d(\kappa)$-matrix. For $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$, denote by $M_{I}(z,\kappa)$ the $d(\kappa)\times d(\kappa)$-minor of that matrix located at the rows with indices in $I$.
@@
6. Orthogonality relations
6.1. Statement
Recall that $d(-\kappa)=\left[\frac{n(p-k)}{p}\right]$.
@@ 6.1. Statement
Let $p>n$ and $10<d(\kappa)<n-$ 1. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
(6.1) $S\big{(}Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big{)}=\sum_{a=1}^{n}Q^{mp-1}_{a}(-z;-\kappa)\,Q^{\ell p-1}_{a}(z;\kappa)=0,$
@@
(6.2) $S\big{(}\bar{Q}^{mp-1}(-z;-\kappa),\bar{Q}^{\ell p-1}(z;\kappa)\big{)}=\sum_{a=1}^{n}\bar{Q}^{mp-1}_{a}(-z;-\kappa)\,\bar{Q}^{\ell p-1}_{a}(z;\kappa).=0.$
Two different proofs of formula (6.2) are given in [VV1, Theorem 3.11] and [VV1, Appendix A].
6.2. Special restrictions
@@
(6.3) $z_{i_{b}}=((b-1)k-1)\kappa,\qquad b=1,\dots,a.$
@@
(6.4) $Q(t,z;\kappa)_{S_{I}}=(t;\kappa)_{ak-1}\,(P_{1}(t,z),\dots,P_{n}(t,z))^{\intercal}$
where $P_{1}(t,z),\dots,P_{n}(t,z)$ are suitable polynomials of degree $(n-a)k$.
@@
(6.6) $G_{\ell,m}(z;\kappa)=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{1\leqslant j_{1}<\dots<j_{b}\leqslant n}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\dots h(z_{j_{b}}),\qquad c_{0},c_{j_{1},\dots,j_{b}}\in{\mathbb{F}}_{p}\,.$
@@
On the one hand, we have $\deg Q^{\ell p-1}(z;\kappa)=kn-\ell p$ and $\deg Q^{mp-1}(-z;-\kappa)=(p-k)n-mp$. Hence $\deg G_{\ell,m}(z;\kappa)\leqslant(n-\ell-m)p$.
@@
The $p$-hypergeometric solution $Q^{\ell p-1}(z,\kappa)$ is associated with the integer $0<k<p$ such that $\kappa k\equiv-1\pmod{p},$ while the $p$-hypergeometric solution $Q^{mp-1}(z,-\kappa)$ is associated with the integer $0<p-k<p$ such that $-\kappa(p-k)\equiv-1\pmod{p}$. Given $m,\ell$,   we say that $k$ is a good parameter if
(6.7) $\ell(p-k)<mk.$
We say that $p-k$ is a good parameter if $mk<\ell(p-k)$. Notice that $mk\neq\ell(p-k)$. Otherwise $p$ must divide $\ell+m$ which is impossible since $\ell+m\leqslant n-1<p$.
Having the two integers $k,\,p-k$ and two solutions $Q^{\ell p-1}(z;\kappa)$ and $Q^{mp-1}(z;-\kappa)$, we may and will assume that $k$ denotes the good parameter.
@@
Let $\{1,\dots,n\}=I\cup J$ be a partition where $|I|=\ell+m$ and $|J|=n-\ell-m$. We may apply formula (6.5) to $Q^{\ell p-1}(z;\kappa)$ with $a=\ell+m$ since $k$ is a good parameter and hence $\ell p<(\ell+m)k$. Hence we have $G_{\ell,m}(z;\kappa)_{I}=0$ where
(6.8) $G_{\ell,m}(z;\kappa)_{I}=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{\{j_{1}<\dots<j_{b}\}\subset J}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\dots h(z_{j_{b}}).$
Hence the right-hand side polynomial at (6.8) is the zero polynomial for any subset $I$ with $|I|=\ell+m$. Therefore, $G_{\ell,m}(z;\kappa)$ is the zero polynomial. ∎
@@
7. Invariant subbundles
7.1. Subbundle of qKZ connection
Assume that $p>n$ and $\kappa\in{\mathbb{F}}_{p}^{\times}$. Consider the discrete qKZ connection on the direct product $V\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$, where
(7.1) $V=\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}$
and the qKZ operators are defined by formula (2.3),
(7.2) $K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\,\dots\,R^{(a,1)}(z_{a}-z_{1}-\kappa)$
×$\hskip 120.0pt\times\ R^{(a,n)}(z_{a}-z_{n})\,\dots\,R^{(a,a+1)}(z_{a}\mskip-0.99998mu-z_{a+1})\,.$

### v2

@@ 4.3 Step $\boldsymbol{-\kappa}$
$Q^{\ell p-1}(z;-\kappa),\qquad\ell=1,\dots,\left[\frac{n(p-k)}{p}\right],$
are solutions of the qKZ equations with step $\kappa$ and values in $V$.
If $p$ does not divide $n$, then the total number of $p$-hypergeometric solutions of the qKZ equations with values in $V$ and steps $\kappa$ and $\kappa$ equals
$\left[\frac{n{k}}{p}\right]+\left[\frac{n(p-{k})}{p}\right]=n-1=\dim V.$
4.4 $\boldsymbol{p}$-hypergeometric solutions of KZ equations
In this subsection, we remind the construction in [6] of polynomial solutions modulo $p$ of the differential KZ equations with values in $V=\operatorname{Sing}L^{\otimes n}[n-2]$.
Let $0<k<p$ be the positive integers such that $\kappa k\equiv-1\pmod{p}$. Define master polynomial $\bar{\Phi}(t,z,\kappa)=\prod_{a=1}^{n}(t-z_{a})^{k}$}. For $1\leqslant a\leqslant n$, define the weight functions
@@ 4.4 $\boldsymbol{p}$-hypergeometric solutions of KZ equations
$\bar{Q}(t,z;\kappa)=\big(\bar{Q}_{1}(t,z;\kappa),\dots,\bar{Q}_{n}(t,z;\kappa)\big)^{\intercal}=\sum_{i}\bar{Q}^{i}(z;\kappa)t^{i},$
where $\bar{Q}^{i}(z;\kappa)=\big(\bar{Q}^{i}_{1}(z;\kappa),\dots,\bar{Q}^{i}_{n}(z;\kappa)\big)^{\intercal}$ are vectors of polynomials in $z$.
Theorem 4.5 ([6]).
@@
$\bar{Q}^{\ell p-1}(z,\kappa),\qquad\ell=1,\dots,\left[\frac{n{k}}{p}\right],$
@@
Notice that $\bar{Q}_{a}(t,z;\kappa)$ are homogeneous polynomials in variables $t$, $z$ of degree nk1nk-1, and $\bar{Q}^{\ell p-1}(z;\kappa)$} are vectors of homogeneous polynomials in $z$ of degree $nk-\ell p$.
4.5 Top-degree part of $\boldsymbol{p}$-hypergeometric solutions
@@
$P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}a_{d_{0},\dots,d_{n+1}}t^{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}}.$ (4.4)
@@
$P(t,z_{1},\dots,z_{n},0)=\sum_{d_{0}+\dots+d_{n}=d}a_{d_{0},\dots,d_{n},0}t^{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}.$ (4.5)
@@
$\bar{P}_{e}(z_{1},\dots,z_{n})=\sum_{e+d_{1}+\dots+d_{n}=d}a_{e,d_{1},\dots,d_{n},0}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}},$
@@
$P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}b_{d_{0},\dots,d_{n+1}}(t;\alpha)_{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}}.$ (4.6)
@@
$P_{e}(z_{1},\dots,z_{n},\alpha)=\sum_{e+d_{1}+\dots+d_{n+1}=d}b_{e,d_{1},\dots,d_{n+1}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}},$
@@
$P_{e}(z_{1},\dots,z_{n},0)=\sum_{e+d_{1}+\dots+d_{n}=d}b_{e,d_{1},\dots,d_{n},0}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}.$
@@
By formula (3.3), we have $a_{e,d_{1},\dots,d_{n},0}=b_{e,d_{1},\dots,d_{n},0}$ for any $e,d_{1},\dots,d_{n}$. This implies the lemma. ∎
Corollary 4.7 ([5]).
Let $Q^{\ell p-1}(z;\kappa)$ be a $p$-hypergeometric solution of the qKZ equations from Theorem 4.3. Then $Q^{\ell p-1}(z;\kappa)=\bar{Q}^{\ell p-1}(z;\kappa)+\cdots$, where $\bar{Q}^{\ell p-1}(z;\kappa)$ is the corresponding $p$-hypergeometric solution of the KZ equations from Theorem 4.5, and the dots denote the terms of degree less than $nk-\ell p=\deg\bar{Q}^{\ell p-1}(z;\kappa)$.
@@
5 Linear independence
5.1 Lexicographical order
Define length-lexicographical order on monomials $v=z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$: $v<w$ if degree(v)<(v)<degree(w)(w) or if degree(v)=(v)=degree(w)(w) and $v$ is smaller than $w$ in lexicographical order.
For a polynomial $f(z)=\sum_{d_{1},\dots,d_{n}}a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$ denote by $\text{L}f(z)$ the nonzero summand $a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$ with the lexicographically largest monomial $z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$. We call $\text{L}f(z)$ the leading term of $f(z)$.
5.2 Leading terms
@@
We have $r(1)>r(2)>\cdots$. Denote
$ga=(n-r(\ell))k-\ell p,\qquad u_{\ell}=\frac{(-1)^{nk-\ell p}}{k}\binom{k}{a}(0,\dots,0,k-a,k,\dots,k)^{\intercal}\in{\mathbb{F}}_{p}^{n},$
where kak-a is the $r(\ell)+1$-st coordinate. Notice that the integers $k$ and kak-a are not divisible by $p$, and $u_{\ell}\in{\mathbb{F}}_{p}^{n}$ is a singular vector since the sum of its coordinates equals zero.
Lemma 5.1 ([7, Lemma 3.1]).
Assume that $p$ does not divide $n$ and $\ell\in\{1,\dots,d(\kappa)\}$. Then the leading term of $\bar{Q}^{\ell p-1}(z;\kappa)$ is given by the formula
$\operatorname{L}\bar{Q}^{\ell p-1}(z;\kappa)=(z_{1}\cdots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}u_{\ell}.$
@@
$\operatorname{L}Q^{\ell p-1}(z;\kappa)=\operatorname{L}\bar{Q}^{\ell p-1}(z;\kappa).$
@@
$\operatorname{L}Q^{\ell p-1}(z;\kappa)=(z_{1}\cdots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}u_{\ell}.$ (5.1)
@@
Let $n=3$ and $d(\kappa)=\big[\frac{3k}{p}\big]=1$. Then $\frac{p}{3}<k<\frac{2p}{3}$. In this case there is exactly one $p$-hypergeometric solution $Q^{p-1}(z;\kappa)$. If $\frac{p}{2}<k<\frac{2p}{3}$, then the leading term of $Q^{p-1}(z;\kappa)$ is
$z_{1}^{k}z_{2}^{2k-p}\frac{(-1)^{3k-p}}{k}\binom{k}{2k-p}(0,p-k,k)^{\intercal}.$
@@
$z_{1}^{3k-p}\frac{(-1)^{k}}{k}\binom{k}{3k-p}(p-2k,k,k)^{\intercal}.$
Consider the collections of the $n$-vectors $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, as an $(n\times d(\kappa))$-matrix. For $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$, denote by $M_{I}(z,\kappa)$ the $(d(\kappa)\times d(\kappa))$-minor of that matrix located at the rows with indices in $I$.
@@
6 Orthogonality relations
6.1 Statement
Recall that $d(-\kappa)=\big[\frac{n(p-k)}{p}\big]$.
@@ 6.1 Statement
Let $p>n$ and $0<d(\kappa)<n-1$. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
$S\big(Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big)=\sum_{a=1}^{n}Q^{mp-1}_{a}(-z;-\kappa)Q^{\ell p-1}_{a}(z;\kappa)=0,$ (6.1)
@@
$S\big(\bar{Q}^{mp-1}(-z;-\kappa),\bar{Q}^{\ell p-1}(z;\kappa)\big)=\sum_{a=1}^{n}\bar{Q}^{mp-1}_{a}(-z;-\kappa)\bar{Q}^{\ell p-1}_{a}(z;\kappa).=0.$ (6.2)
Two different proofs of formula (6.2) are given in [8, Theorem 3.11] and [8, Appendix A].
6.2 Special restrictions
@@
$z_{i_{b}}=((b-1)k-1)\kappa,\qquad b=1,\dots,a.$ (6.3)
@@
$Q(t,z;\kappa)_{S_{I}}=(t;\kappa)_{ak-1}(P_{1}(t,z),\dots,P_{n}(t,z))^{\intercal},$
where $P_{1}(t,z),\dots,P_{n}(t,z)$ are suitable polynomials of degree (na)k(n-a)k.
@@
$G_{\ell,m}(z;\kappa)=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{1\leqslant j_{1}<\dots<j_{b}\leqslant n}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\cdots h(z_{j_{b}}),\qquad c_{0},c_{j_{1},\dots,j_{b}}\in{\mathbb{F}}_{p}.$
@@
On the one hand, we have
$\deg Q^{\ell p-1}(z;\kappa)=kn-\ell p\qquad\text{and}\qquad\deg Q^{mp-1}(-z;-\kappa)=(p-k)n-mp.$
Hence $\deg G_{\ell,m}(z;\kappa)\leqslant(n-\ell-m)p$.
@@
The $p$-hypergeometric solution $Q^{\ell p-1}(z,\kappa)$ is associated with the integer $0<k<p$ such that $\kappa k\equiv-1\pmod{p}$}, while the $p$-hypergeometric solution $Q^{mp-1}(z,-\kappa)$ is associated with the integer 0<pk<p0<p-k<p such that $\kappa(p-k)\equiv-1\pmod{p}$. Given $m$, $\ell$, we say that $k$ is a good parameter if $\ell(p-k)<mk$. We say that pkp-k is a good parameter if $mk<\ell(p-k)$. Notice that $mk\neq\ell(p-k)$. Otherwise $p$ must divide $\ell+m$ which is impossible since $\ell+m\leqslant n-1<p$.
Having the two integers $k$, pkp-k and two solutions $Q^{\ell p-1}(z;\kappa)$ and $Q^{mp-1}(z;-\kappa)$, we may and will assume that $k$ denotes the good parameter.
@@
Let $\{1,\dots,n\}=I\cup J$ be a partition where $|I|=\ell+m$ and $|J|=n-\ell-m$. We may apply formula (6.4) to $Q^{\ell p-1}(z;\kappa)$ with $a=\ell+m$ since $k$ is a good parameter and hence $\ell p<(\ell+m)k$. Hence we have $G_{\ell,m}(z;\kappa)_{I}=0$, where
$G_{\ell,m}(z;\kappa)_{I}=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{\{j_{1}<\dots<j_{b}\}\subset J}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\dots h(z_{j_{b}}).$ (6.5)
Hence the right-hand side polynomial at (6.5) is the zero polynomial for any subset II with $|I|=\ell+m$. Therefore, $G_{\ell,m}(z;\kappa)$ is the zero polynomial. ∎
@@
7 Invariant subbundles
7.1 Subbundle of qKZ connection
Assume that $p>n$ and $\kappa\in{\mathbb{F}}_{p}^{\times}$. Consider the discrete qKZ connection on the direct product $V\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$}, where $V=\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}$ and the qKZ operators are defined by formula (2.2),
$K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\cdots R^{(a,1)}(z_{a}-z_{1}-\kappa)$
[…]